OpenStudy (anonymous):

solve using the sqrt property of equality: x^2-6x+9=50

OpenStudy (anonymous):

(x-3)^2=50 x=3+/- radical 50

OpenStudy (anonymous):

\[x = 3 - 5\sqrt{2}\] and \[x = 3 + 5\sqrt{2}\]